# Omni-R1: Do You Really Need Audio to Fine-Tune Your Audio LLM?

### v1

@@ Abstract
We propose Omni-R1 which fine-tunes a recent multi-modal LLM, Qwen2.5-Omni, on an audio question answering dataset with the reinforcement learning method GRPO. This leads to new State-of-the-Art performance on the recent MMAU benchmark. Omni-R1 achieves the highest accuracies on the sounds, music, speech, and overall average categories, both on the Test-mini and Test-full splits. To understand the performance improvement, we tested models both with and without audio and found that much of the performance improvement from GRPO could be attributed to better text-based reasoning. We also made a surprising discovery that fine-tuning without audio on a text-only dataset was effective at improving the audio-based performance.
@@ I Introduction
Reinforcement Learning (RL) has recently been shown to improve the reasoning capabilities of Large Language Models (LLMs) [1]. We are motivated by these advancements to improve the capabilities of Audio LLMs - models which take in audio input and text and can perform tasks such as question answering. Building on Qwen2.5-Omni [2], a State-of-the-Art (SOTA) multi-modal LLM, we introduce Omni-R1 using a fine-tuning pipeline based on the RL method Group Relative Policy Optimization (GRPO) [3]. Using a simple prompt, our model forgoes complex chain-of-thought or structured reasoning outputs and instead directly outputs answer choices. We use the recent MMAU [4] benchmark to test our models. Fine-tuning on the audio and human-annotated questions from the AVQA [5] dataset boosts Qwen2.5-Omni’s average accuracy on MMAU Test-mini from 65.9% to 68.6%, and on the Test-full split from 68.4% to 70.8%.
We also propose to automatically generate audio question answering datasets. We prompt ChatGPT with several audio captions from an audio LLM to generate questions and answer choices for the 40k audios in AVQA and 182k audios in VGGSound. Scaling up the training data with our automatically generated questions results in further gains, achieving new SOTA of 71.3% on Test-mini and 71.2% on Test full.
@@
  • Omni-R1, a streamlined GRPO fine-tuning on Qwen2.5-Omni that achieves new SOTA on MMAU without any complex prompts or explicit reasoning.
@@ II Related Work
The MMAU benchmark [4] is a recent large-scale audio question answering dataset designed to test audio LLMs. It is a fixed answer choice dataset providing one correct answer and three incorrect answers for each question. It contains questions about sounds, speech, and music of varying difficulty. Some of the questions require external world-based knowledge that is not provided in the questions themselves.
GRPO is a Reinforcement Learning (RL) method proposed for instilling better reasoning capabilities in LLMs [3, 1]. Towards better Audio LLMs, R1-AQA [6] proposes to use GRPO by fine-tuning Qwen2-Audio [7] on the AVQA [5] dataset of audio question and answers. This method achieved the previous SOTA on MMAU, and our approach is inspired by them. Using GRPO, we fine-tune Qwen2.5-Omni-7B [2], a recent multi-modal LLM proposed to handle both audio and video inputs, on AVQA. Doing so, we were able to achieve the new SOTA on MMAU. Further, we propose to automatically generate audio question and answer training data, which leads to even better performance.
@@ III-A Audio LLM and Prompt
We propose Omni-R1 which fine-tunes Qwen2.5-Omni-7B [2] using GRPO. During fine-tuning and inference, we use the prompt: <question> Please choose the answer from the following options: <choices>. We chose this prompt since it is simple and the model learns to directly output one of the answer choices. This is especially useful since we have limited GPU memory. Thanks to this simple prompt, we were able to perform full-finetuning on GPUs with only 48GB GPU memory, in contrast to recent works which perform fine-tuning on GPUs with 80GB GPU memory [6]. Although recent works [10, 6, 8] propose more complex prompts with reasoning, our simple prompt was sufficient for our models to achieve the new SOTA on MMAU.
@@ III-B Reinforcement Learning
TABLE I: Accuracies (%) on MMAU. SARI calculates scores using an LLM as a judge instead of the usual string matching.
@@ III-C Automatic Q/A Dataset Generation with ChatGPT
We propose to generate audio question answer datasets automatically. Given an audio dataset, we first acquire audio-based text captions from an audio LLM. Specifically, we use the captions generated by Qwen-2 Audio [7] in the AudioSetCaps dataset [17], which is a compilation of captions generated for many open-source audio datasets. We use multiple captions which describe the overall acoustic scene, the speech present in the audio, and the music present in the audio. Next, we use the captions and prompt ChatGPT to generate questions and answers using the provided captions. We instruct the model to generate a correct answer and three other answers which are similar but clearly incorrect. We reviewed the generated questions and answers and found the quality to be reasonable, however, we noticed that the original captions feature hallucinations especially about music not present in the audio. The generated questions could then ask about some music or sounds not actually present in the audio. We also manually randomized the order of the correct answer within the four choices to ensure a balance.
@@ IV-A Experimental Setup
IV-B Main Results
Table I shows the main results on MMAU [4]. For the baselines, we show the recent methods which achieve State-of-the-Art (SOTA) performance. We refer readers to MMAU [4] for the older baselines [19, 20, 21, 22, 23, 24, 25, 7, 26, 27]. Note that all of the methods were evaluated on the Test-mini split with 1k audio samples, while only a few of them evaluated on the Test-full split with 9k audio samples.
@@ IV-B Main Results
TABLE II: Ablation of RL fine-tuning with vs. without audio across datasets. Accuracies (%) on MMAU mini at inference time.
@@
IV-C Fine-Tuning Audio LLMs without Audio
In this section, we investigate how the models improve with GRPO fine-tuning. In particular, we wanted to understand why GRPO improved Qwen2.5-Omni’s performance less than Qwen2.5-Omni’s. Comparing the result of fine-tuning Qwen2.5-Omni with GRPO (Omni-R1, ours) and Qwen2-Audio with GRPO (R1-AQA[6]) on AVQA, Qwen2.5-Omni’s average score improved from 65.9% to 68.6% (2.7% absolute) while Qwen2-Audio’s average score improved from 52.7% to 65.6% (12.9% absolute). To investigate this, we tested the models without audio to understand their text reasoning capabilities. We then fine-tune the models without audio (just text) and measure the resulting performance.
Table II shows the result of performing inference both with audio and without audio. The scores reported are the averages on MMAU Test-mini split. For inference without audio, we simply drop the audio tokens from the input. Without access to the input audio, Qwen2-Audio and Qwen2.5-Omni achieve 30.5% and 49.3% respectively. While the performance for Qwen2-Audio is around chance (25%), it is surprising that Qwen2.5-Omni performs so well without audio. It shows that many of the MMAU questions can be answered without audio and that Qwen2.5-Omni has good text-based knowledge about audio. We suspected that GRPO helped Qwen2-Audio more significantly because its text-based reasoning was worse than Qwen2.5-Omni’s. Therefore, we checked the performance after fine-tuning both models with GRPO on AVQA. Qwen2-Audio’s performance without audio increased significantly from 30.5% to 44.6%, while Qwen2.5-Omni’s performance without audio increased only slightly from 49.3% to 51.7%. This shows that fine-tuning with GRPO significantly improved Qwen2-Audio’s text-based reasoning, but not Qwen2.5-Omni since the improvement was smaller. This is reasonable since Qwen2.5-Omni’s performance without audio was already high.
@@ V Conclusion
We propose Omni-R1, an Audio LLM based on fine-tuning the multi-modal LLM Qwen2.5-Omni with GRPO for better audio question answering. We use a simple yet effective prompt which is straight to the point and makes training and testing efficient. On the MMAU benchmark, Omni-R1 sets new State-of-the-Art results across sounds, music, speech, and overall accuracy on both Test-mini and Test-full. To scale up, we also introduced two large, auto-generated audio question answering datasets (AVQA-GPT and VGGS-GPT), which further boost performance. Through our experiments of testing the models with and without audio, we showed that much of the gains come from better text-based reasoning. Surprisingly, fine-tuning with only text (no audio) also delivers strong improvements in audio question answering.

### v2

@@ Abstract
We propose Omni-R1 which fine-tunes a recent multi-modal LLM, Qwen2.5-Omni, on an audio question answering dataset with the reinforcement learning method GRPO. This leads to new State-of-the-Art performance on the recent MMAU and MMAR benchmarks. On MMAU, Omni-R1 achieves the highest accuracies on the sounds, music, speech, and overall average categories, both on the Test-mini and Test-full splits. To understand the performance improvement, we tested models both with and without audio and found that much of the performance improvement from GRPO could be attributed to better text-based reasoning. We also made a surprising discovery that fine-tuning without audio on a text-only dataset was effective at improving the audio-based performance.
@@ I Introduction
Reinforcement Learning (RL) has recently been shown to improve the reasoning capabilities of Large Language Models (LLMs) [1]. We are motivated by these advancements to improve the capabilities of Audio LLMs - models which take in audio input and text and can perform tasks such as question answering. Building on Qwen2.5-Omni [2], a State-of-the-Art (SOTA) multi-modal LLM, we introduce Omni-R1 using a fine-tuning pipeline based on the RL method Group Relative Policy Optimization (GRPO) [3]. Using a simple prompt, our model forgoes complex chain-of-thought or structured reasoning outputs and instead directly outputs answer choices. We use the recent MMAU [4] and MMAR [5] benchmarks to test our models. Fine-tuning on the audio and human-annotated questions from the AVQA [6] dataset boosts Qwen2.5-Omni’s average accuracy on MMAU Test-mini from 65.9% to 68.6%, and on the Test-full split from 68.4% to 70.8%.
We also propose to automatically generate audio question answering datasets. We prompt ChatGPT with several audio captions from an audio LLM to generate questions and answer choices for the 40k audios in AVQA and 182k audios in VGGSound. Scaling up the training data with our automatically generated questions results in further gains, achieving new SOTA of 71.3% on Test-mini and 71.2% on Test full. We obtained similar gains on the MMAR [5] benchmark leading to SOTA results.
@@
  • Omni-R1, a streamlined GRPO fine-tuning on Qwen2.5-Omni that achieves new SOTA on MMAU and MMAR without any complex prompts or explicit reasoning.
@@ II Related Work
Several recent large-scale audio question answering datasets were proposed to test audio LLMs, including MMAU [4] and MMAR [5]. The MMAU benchmark [4] is a fixed answer choice dataset providing one correct answer and three incorrect answers for each question. It contains questions about sounds, speech, and music of varying difficulty. Some of the questions require external world-based knowledge that is not provided in the questions themselves. MMAR [5] is a more recent benchmark designed to elicit deep reasoning from audio LLMs. It includes questions about sounds, speech, and music, and also mixtures of them (for example, audio natively containing both sounds and speech). The question and answer format is the same as MMAU.
GRPO is a Reinforcement Learning (RL) method proposed for instilling better reasoning capabilities in LLMs [3, 1]. Towards better Audio LLMs, R1-AQA [7] proposes to use GRPO by fine-tuning Qwen2-Audio [8] on the AVQA [6] dataset of audio question and answers. This method achieved the previous SOTA on MMAU, and our approach is inspired by them. Using GRPO, we fine-tune Qwen2.5-Omni-7B [2], a recent multi-modal LLM proposed to handle both audio and video inputs, on AVQA. Doing so, we were able to achieve the new SOTA on MMAU and also demonstrate strong performance on the MMAR benchmark, achieving SOTA results among open-source models and competitive performance to significantly bigger, close-ended models. Further, we propose to automatically generate audio question and answer training data, which leads to even better performance.
@@
Figure 1: The prompt used with ChatGPT to generate our AVQA-GPT and VGGS-GPT datasets (Section III-C).
@@ III-A Audio LLM and Prompt
We propose Omni-R1 which fine-tunes Qwen2.5-Omni-7B [2] using GRPO. During fine-tuning and inference, we use the prompt: <question> Please choose the answer from the following options: <choices>. We chose this prompt since it is simple and the model learns to directly output one of the answer choices. This is especially useful since we have limited GPU memory. Thanks to this simple prompt, we were able to perform full-finetuning on GPUs with only 48GB GPU memory, in contrast to recent works which perform fine-tuning on GPUs with 80GB GPU memory [7]. Although recent works [11, 7, 9] propose more complex prompts with reasoning, our simple prompt was sufficient for our models to achieve the new SOTA on MMAU and MMAR.
@@ III-B Reinforcement Learning
TABLE I: Accuracies (%) on MMAU (original). SARI calculates scores using an LLM as a judge instead of the usual string matching.
@@ III-C Automatic Q/A Dataset Generation with ChatGPT
We propose to generate audio question answer datasets automatically. Given an audio dataset, we first acquire audio-based text captions from an audio LLM. Specifically, we use the captions generated by Qwen-2 Audio [8] in the AudioSetCaps dataset [18], which is a compilation of captions generated for many open-source audio datasets. We use multiple captions which describe the overall acoustic scene, the speech present in the audio, and the music present in the audio. Next, we use the captions and prompt ChatGPT to generate questions and answers using the provided captions. We instruct the model to generate a correct answer and three other answers which are similar but clearly incorrect (prompt shown in Figure 1). We reviewed the generated questions and answers and found the quality to be reasonable, however, we noticed that the original captions feature hallucinations especially about music not present in the audio. The generated questions could then ask about some music or sounds not actually present in the audio. We also manually randomized the order of the correct answer within the four choices to ensure a balance.
@@
TABLE II: Accuracies (%) on MMAU (v05.15.25). Qwen2.5 scores are from https://sakshi113.github.io/mmau_homepage/
Model Method MMAU Test-mini (1k audios) MMAU Test-full (9k audios)
Sound Music Speech Avg. Sound Music Speech Avg.
Baselines:
Qwen2.5-Omni-7B Direct Inference [2] - - - - 74.7 68.6 76.0 73.1
Qwen2.5-Omni-7B Reproduced 78.1 65.9 70.6 71.5 74.9 68.0 72.4 71.8
Ours:
Qwen2.5-Omni-7B Omni-R1 (AVQA) 82.0 70.1 75.4 75.8 77.8 69.1 76.3 74.4
Qwen2.5-Omni-7B Omni-R1 (AVQA-GPT) 82.3 72.8 75.4 76.8 78.9 70.9 76.0 75.3
Qwen2.5-Omni-7B Omni-R1 (VGGS-GPT) 81.7 73.4 76.0 77.0 78.3 70.8 75.8 75.0
@@ IV Experiments
In this section, we report the results on the MMAU [4] and MMAR benchmarks. There are two versions of MMAU. The first is the original version, which we denote as MMAU (Original). All of the previous and concurrent methods have been evaluated on this version. Very recently, an updated version of the benchmark was proposed [^1] https://github.com/Sakshi113/MMAU, which we denote as MMAU (v05.15.25). According to the authors, 25% of the questions and answers have been revised for improved clarity and quality, while 5% of the audio files have been refined to enhance consistency and fidelity. We present the results on MMAU (Original) in Section IV-B and on MMAU (v05.15.25) in Section IV-C. We present the results on MMAR in Section IV-D.
@@ IV-A Experimental Setup
IV-B Main Results on MMAU (Original)
Table I shows the main results on MMAU (Original) [4]. For the baselines, we show the recent methods which achieve State-of-the-Art (SOTA) performance. We refer readers to MMAU [4] for the older baselines [20, 21, 22, 23, 24, 25, 26, 8, 27, 28]. Note that all of the methods were evaluated on the Test-mini split with 1k audio samples, while only a few of them evaluated on the Test-full split with 9k audio samples.
@@ IV-B Main Results on MMAU (Original)
TABLE III: Accuracies (%) on MMAR. Mix1: Sound-Music; Mix2: Sound-Speech; Mix3: Music-Speech; Mix4: Sound-Music-Speech.
Model Method Single Modality Mixed Modalities Avg (%)
Sound Music Speech Mix1 Mix2 Mix3 Mix4
Closed-Source Models:
Gemini 2.0 Flash Direct Inference [29] 61.2 51.0 72.1 81.8 72.5 65.9 70.8 65.6
GPT-4o Audio Direct Inference [30] 53.9 51.0 70.4 63.6 72.5 62.2 75.0 63.5
Open-Source Models:
[…]
Ours:
[…]
IV-C Main Results on MMAU (v05.15.25)
Table II shows the results on MMAU (v05.15.25), the updated version of the benchmark. We focus on Qwen2.5-Omni-7B since it achieves the highest scores on this version. The authors of MMAU reported the scores of Qwen2.5-Omni-7B on the benchmark website with the Test-full split but not with the Test-mini split. We evaluated Qwen2.5-Omni-7B on both splits and the model achieved average scores of 71.5% on Test-mini and 71.8% on Test-full. Comparing our reproduced results vs the results presented on the benchmark website using Test-full, our reproduced results are slightly lower (71.8% reproduced vs 73.1% reported). Since the Test-full answers are hidden and the scores are computed by an online server, we have reached out to the authors to inquire why our reproduced results are slightly worse than theirs.
Next, we show our Omni-R1 models which fine-tune Qwen2.5-Omni with GRPO. Similar to the results on MMAU (Original), Omni-R1 achieves significant improvements over the base model on MMAU (v05.15.25). First, fine-tuning on AVQA results in average scores of 75.8% and 74.4% on Test-mini and Test-full respectively. This shows the effectiveness of the GRPO fine-tuning process. Second, fine-tuning on AVQA-GPT results in average scores of 76.8% and 75.3% on Test-mini and Test-full respectively. Fine-tuning on AVQA-GPT results in better scores than fine-tuning on AVQA, which shows the effectiveness of our automatically generated question and answers. The average score of 75.3% is our best result on Test-Full and it outperforms Qwen2.5-Omni-7B’s average score of 71.8% on Test-full by 3.5% absolute. Finally, fine-tuning on AVQA-GPT results in average scores of 77.0% and 75.0% on Test-mini and Test-full respectively. The average score of 77.0% is our best result on Test-mini and it outperforms Qwen2.5-Omni-7B’s average score of 71.5% on Test-full by 5.5% absolute. Overall these results show that our fine-tuning process is robust and shows consistent gains compared to the baseline method even with updates to the benchmark.
IV-D Main Results on MMAR
Table III shows the results on MMAR [5], which only provides a test set of 1k audios. For the baselines, we show the results of Qwen2-Audio and Qwen-2.5-Omni and refer the reader to the original paper for the other baselines. Our reproduction of the base Qwen-2.5-Omni results in an average score of 58%.
Next, we show our Omni-R1 models which fine-tune Qwen2.5-Omni with GRPO. First, Omni-R1 fine-tuned on AVQA achieves an average score of 61.2% (3.2% absolute higher than the base model), showing the effectiveness of our fine-tuning method. Next, fine-tuning on our automtically generated AVQA-GPT and VGGS-GPT datasets results in further gains. Our best model fine-tuned on VGGS-GPT achieves an average score of 63.4% (5.4% absolute higher than the base model), setting a new State-of-the-Art for open-source models. Our fine-tuned models improved the scores on all categories compared to the base model, except on mixtures of sound and music (Mix1). However, this subset only contains 11 test samples in total and the number of correct answers changed from 6 to 5 which is acceptable. Our model’s performance approaches the scores of strong closed-source models such as Gemini and GPT-4o, shown at the top of the table. Finally, we note that while the MMAR dataset was designed to elicit deep reasoning from audio LLMs, our model could achieve new State-of-the-Art results without outputting long text reasoning steps and simply outputting the correct answer.
IV-E Fine-Tuning Audio LLMs without Audio
In this section, we investigate how the models improve with GRPO fine-tuning. In particular, we wanted to understand why GRPO improved Qwen2.5-Omni’s performance less than Qwen2.5-Omni’s on MMAU (Original). Comparing the result of fine-tuning Qwen2.5-Omni with GRPO (Omni-R1, ours) and Qwen2-Audio with GRPO (R1-AQA[7]) on AVQA in Table I, Qwen2.5-Omni’s average score improved from 65.9% to 68.6% (2.7% absolute) while Qwen2-Audio’s average score improved from 52.7% to 65.6% (12.9% absolute). To investigate this, we tested the models without audio to understand their text reasoning capabilities. We then fine-tune the models without audio (just text) and measure the resulting performance. Note that we use MMAU (Original) for these experiments.
Table IV shows the result of performing inference both with audio and without audio. The scores reported are the averages on MMAU Test-mini split. For inference without audio, we simply drop the audio tokens from the input. Without access to the input audio, Qwen2-Audio and Qwen2.5-Omni achieve 30.5% and 49.3% respectively. While the performance for Qwen2-Audio is around chance (25%), it is surprising that Qwen2.5-Omni performs so well without audio. It shows that many of the MMAU questions can be answered without audio and that Qwen2.5-Omni has good text-based knowledge about audio. We suspected that GRPO helped Qwen2-Audio more significantly because its text-based reasoning was worse than Qwen2.5-Omni’s. Therefore, we checked the performance after fine-tuning both models with GRPO on AVQA. Qwen2-Audio’s performance without audio increased significantly from 30.5% to 44.6%, while Qwen2.5-Omni’s performance without audio increased only slightly from 49.3% to 51.7%. This shows that fine-tuning with GRPO significantly improved Qwen2-Audio’s text-based reasoning, but not Qwen2.5-Omni since the improvement was smaller. This is reasonable since Qwen2.5-Omni’s performance without audio was already high.
@@ IV-E Fine-Tuning Audio LLMs without Audio
TABLE IV: Ablation of RL fine-tuning with vs. without audio across datasets. Accuracies (%) on MMAU mini at inference time.
@@ V Conclusion
We propose Omni-R1, an Audio LLM based on fine-tuning the multi-modal LLM Qwen2.5-Omni with GRPO for better audio question answering. We use a simple yet effective prompt which is straight to the point and makes training and testing efficient. On the MMAU benchmark, Omni-R1 sets new State-of-the-Art results across sounds, music, speech, and overall accuracy on both Test-mini and Test-full. On the MMAR benchmark, our method achieves new State-of-the-Art results among open-source models and competitive performance compared to significantly larger, closed-source models. To scale up, we also introduced two large, auto-generated audio question answering datasets (AVQA-GPT and VGGS-GPT), which further boost performance. Through our experiments of testing the models with and without audio, we showed that much of the gains come from better text-based reasoning. Surprisingly, fine-tuning with only text (no audio) also delivers strong improvements in audio question answering.